Two parallel lines are intersected be a transversal forming alternate interior angles whose measures in degrees are described as : 7x-1 and 3x+7. What are their degrees in measures?
Answer by richwmiller(17219)

7x-1 = 3x+7
4x=8
x=2
3x+7=13
7x-1=13